Circular No. 7137

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999bg IN IC 758 K. Ayani, Bisei Astronomical Observatory (BAO); and H. Yamaoka,

Kyushu University, report: “We have obtained a CCD spectrum (range 400–750 nm) of SN 1999bg (cf. IAUC 7135) on Mar. 31.6 UT with the 1.01m BAO telescope. Preliminary reduction reveals that, although affected by moonlight, the spectrum shows an almost-featureless continuum with a weak, broad absorption feature at ∼ 470 nm. There probably exists a weak P Cyg-like feature in the range 625–680 nm. If these belong to Hβ and Hα, so that the supernova is of type II, the expansion velocity would be ∼ 11 000 km/s. The narrow Hα emission, likely due to the surrounding H II region at the position on a Digitized Sky Survey image, shows that the recession velocity of IC 758 is ∼ 1100 km/s.”

COMET C/1998 M5 (LINEAR) Ephemeris extension to IAUC 7115 (orbital elements MPC 34126):

1999TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Apr. 2 8^{h}04.^{m}4 +72^{o}05.^{′}7 1.642 1.947 91.^{o}6 30.^{o}9 10.0
7 8 09.39 +67 20.9 1.698 1.976 90.3 30.4 10.1
12 8 14.28 +62 51.2 1.763 2.006 88.6 30.0 10.3
17 8 19.11 +58 37.9 1.836 2.038 86.5 29.4 10.4
22 8 23.92 +54 41.6 1.918 2.071 84.1 28.9 10.6
27 8 28.72 +51 02.2 2.005 2.105 81.4 28.2 10.7
May 2 8 33.50 +47 39.0 2.098 2.140 78.6 27.5 10.9
7 8 38.26 +44 30.9 2.195 2.176 75.6 26.7 11.1
12 8 43.02 +41 36.8 2.296 2.214 72.6 25.8 11.3
17 8 47.78 +38 55.3 2.399 2.252 69.4 24.9 11.4
22 8 52.52 +36 25.3 2.503 2.290 66.2 23.8 11.6
27 8 57.26 +34 05.6 2.609 2.330 62.9 22.8 11.8
June 1 9 01.97 +31 55.1 2.715 2.370 59.7 21.7 11.9
6 9 06.66 +29 52.7 2.820 2.411 56.4 20.5 12.1
11 9 11.33 +27 57.6 2.925 2.453 53.0 19.3 12.2
16 9 15.98 +26 08.8 3.028 2.494 49.7 18.1 12.4
21 9 20.61 +24 25.7 3.129 2.537 46.3 16.8 12.5

Total visual magnitude estimates: Mar. 12.83 UT, 8.7 (K. Hornoch, Lelekovice, Czech Republic, 10×80 binoculars); 17.03, 9.1 (J. E. Bortle, Stormville, NY, 20×80 binoculars); 22.80, 8.9 (M. Reszelski, Szamotuly, Poland, 20×60 binoculars); 27.12, 8.8 (Hornoch).

1999 April 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green